OpenStudy (anonymous):

A single die is rolled. Find the probability of rolling an even number or a number less than 6. Is it 5/6?

OpenStudy (anonymous):

Even numbers on a dice: 2,4,6 = 1/2 chance. Numbers less than 6 on a dice: 5/6 chance. 1/2 + 5/6 = 8/6. 8/6 is impossible so we'll take it down to 1. There will ALWAYS be an even number, or a number less than 6. The answer is 1.

OpenStudy (anonymous):

Ohh, I get it. Thank you so much!